OpenStudy (anonymous):

I have an aswer can someone check me please...... Find the derivative of b(x)= ((e^x)/(2x+3))

OpenStudy (anonymous):

\[\frac{ e^x }{ 2x+3 }\]

OpenStudy (anonymous):

wait minute

OpenStudy (anonymous):

My answer that I got was\[\frac{ xe^x-2e^x }{ (2x+3)^2}\]

OpenStudy (phi):

you can use wolfram to check your result but it is incorrect.

OpenStudy (anonymous):

how is it incorrect

OpenStudy (phi):

how did you get it ? I would write the problem as e^x (2x+3)^(-1) and use the product rule

OpenStudy (anonymous):

\[\frac{ e ^{x}(2x+3)-2e ^{x} }{ (2x-3)^{2} }\]

OpenStudy (anonymous):

ok I think I got it would the answer be \[b(x)=\frac{ -2xe^x+e^x }{ (2x+3)^2 }\]

OpenStudy (anonymous):

the rule is let Primarily=P ,Numerator=P , derivative=d \[\frac{ (dN * P) - (dP * N)}{ P ^{2} }\]

OpenStudy (phi):

closer. here is one way to do it d/dx ( e^x (2x+3)^(-1) )= e^x d/dx (2x+3)^-1 + (2x+3)^-1 d/dx e^x the d/dx (2x+3)^-1 = -(2x+3)^-2 * 2= -2/(2x+3)^2 and the first term is -2e^x/(2x+3)^2 the 2nd term is e^x/(2x+3)= e^x(2x+3)/(2x+3)^2 \[ \frac{-2e^x + e^x(2x+3)}{(2x+3)^2} \] or \[ \frac{3e^x -2e^x + 2xe^x}{(2x+3)^2} \] \[ \frac{e^x + 2xe^x}{(2x+3)^2} \]

OpenStudy (anonymous):

ok but would this work\[b(x)=\frac{ 2xe^x+e^x }{ (2x+3)^2 }\]

OpenStudy (anonymous):

sory Numerator=N in my answer

OpenStudy (phi):

yes, that is the same thing

OpenStudy (anonymous):

awesome thanks for your help

OpenStudy (anonymous):

right